### v1

@@ 6.1 Static Equilibrium at Local Maxima and Minima
Our formulation based on the kinetic energy is akin to the force-norm minimization [Twigg:2011:OSS:2019406.2019437] (which is also related to the nonlinear force solves with inverse dynamics [Hadap2006strands, Featherstone2016], ANM [Chen2014asymptotic, Jia2021SANM], and global-local approaches [Hsu2022sag, Hsu2023sag]) and is designed to find a static equilibrium where forces are zero. While the achieved static equilibrium is at a local minimum for the rest shape optimization (37), it can be at a local maximum for the optimization on forward simulation (1), as discussed in [Derouet-Jourdan2010, Hsu2022sag]. As such, some perturbations on the generalized positions can relatively easily break the static equilibrium; however, this case is rare for strands with one-sided clamping since such strands typically do not experience compression and settle at a local minimum for (1) via the rest shape optimization (37). If one needs to guarantee a static equilibrium at a local minimum with respect to both generalized positions and rest shape parameters, it seems promising to use local material stiffening [Derouet-Jourdan2010, Hsu2022sag] and the adjoint-based approaches [Perez2015rod, thuerey2021pbdl].
@@ 6.2 Material Stiffness Parameter Optimization
While our rest shape optimizer was able to find rest shape parameters that achieve the desired static equilibrium in various settings with reasonable material stiffness parameters, in general, it is not guaranteed that such parameters exist, e.g., when the given system is overdetermined or such parameters are outside of the box constraints. In addition, when the rest shape is significantly modified to achieve the static equilibrium (even under the box constraints, which are designed to avoid introducing stability problems), the strand shape may not return to the original configuration after the strand moves dynamically because it can get stuck in different local extrema (1). In these cases, increasing the material stiffness can reduce the required rest shape changes and can be helpful to establish a stabler static equilibrium (albeit compromising user-desired elastic dynamics)[Derouet-Jourdan2010, Twigg:2011:OSS:2019406.2019437, Hsu2022sag, Hsu2023sag].
@@ 6.3 Optimization Solvers
We designed our optimization solvers by seeking a balance between performance and simplicity, but it is possible that other alternatives could potentially perform better. Although the penalty method was quite satisfactory in our examples (since small violations of box constraints or clamping were permissible), other constrained optimization solvers (e.g., augmented Lagrangian method (ALM) [NoceWrig06, Takahashi2021], interior point method (IPM) [Takahashi2024qp], and active-set method [Dostal2005, Takahashi2023]) could be employed to more accurately satisfy the box constraints. Other nonlinear solvers (e.g., nonlinear conjugate gradient, quasi-Newton, and LMA [NoceWrig06]) are also available, and it would be worth comparing their performance with Gauss-Newton. While we arranged the rest shape parameters in their current order (rest length, rest curvature, rest twist) so that the rest twist can easily be added and removed from the rest shape optimization, by interleaving these parameters (as we did for the generalized positions), the inner linear systems for Newton-type optimizers would possess a banded structure that enables more efficient direct solves without reordering of the system(s).
@@ 6.4 Toward More General Inverse Problems
In our framework, for simplicity and efficiency, we assumed a strand with constant radius, density, and material stiffness. It would be worthwhile removing this assumption to support more general anisotropic and inhomogeneous strands [Hafner2021curves, Hafner2023rod]. Moreover, our framework should be able to extend to support two-dimensional shells and three-dimensional volumetric structures, along with more rigorous analysis for the box constraints imposed on the rest shape parameters. In addition, although we also assumed no contacts among strands or other objects, this approach can lead to larger rest shape changes than necessary in cases where the elastic strands could be supported by each other or other objects when the simulation starts [Derouet-Jourdan:2013:IDH:2508363.2508398, Ly2018shell, Hsu2022sag, Hsu2023sag]. Thus, a promising extension would be to take frictional contacts into account during the rest shape optimization.

### v2

@@ 6.1 Static Equilibrium at Local Maxima and Minima
Our formulation based on the kinetic energy is akin to the force-norm minimization [TKA11] (which is also related to the nonlinear force solves with inverse dynamics [Had06, Fea16], ANM [CZXZ14, Jia21], and global-local approaches [HTYW22, HWP23]) and is designed to find a static equilibrium where forces are zero. While the achieved static equilibrium is at a local minimum for the rest shape optimization (37), it can be at a local maximum for the optimization on forward simulation (1) [DJBDT10, HTYW22]. As such, perturbations on the generalized positions can relatively easily break the static equilibrium; however, this case is rare for strands with one-sided clamping since such strands typically do not experience compression and settle at a local minimum for (1) via the rest shape optimization (37). If one needs to guarantee a static equilibrium at a local minimum with respect to both generalized positions and rest shape parameters, it seems promising to use local material stiffening [DJBDT10, HTYW22] and the adjoint-based approaches [PTC15].
@@ 6.2 Material Stiffness Parameter Optimization
While our rest shape optimizer was able to find rest shape parameters that achieve the desired static equilibrium in various settings with reasonable material stiffness parameters, in general, it is not guaranteed that such parameters exist, e.g., when the given system is overdetermined or such parameters are outside of the box constraints (see Figures 4 and 8). In addition, when the rest shape is significantly modified to achieve the static equilibrium (even under the box constraints, which are designed to avoid introducing stability problems), the strand shape may not return to the original configuration after the strand moves dynamically because it can get stuck in different local extrema (1). In these cases, increasing the material stiffness (or equivalently increasing the strand radius) can reduce the required rest shape changes and can be helpful to establish a stabler static equilibrium (albeit compromising user-desired elastic dynamics)[DJBDT10, TKA11, HTYW22, HWP23].
@@ 6.3 Optimization Solvers
We designed our optimization solvers by seeking a balance between performance and simplicity, but it is possible that other alternatives could potentially perform better. Although the penalty method was quite satisfactory in our examples (since small violations of box constraints or clamping were permissible), other constrained optimization solvers (e.g., augmented Lagrangian method (ALM) [NW06, TB21], interior point method (IPM) [TB24], and active-set method [DS05, TB23]) could be employed to more accurately satisfy the box constraints. Other nonlinear solvers (e.g., nonlinear conjugate gradient, quasi-Newton, and LMA [NW06]) are also available, and it would be worth comparing their performance with Gauss-Newton. While we arranged the rest shape parameters in their current order (rest length, rest curvature, rest twist) so that the rest twist can easily be added and removed from the rest shape optimization, by interleaving these parameters (as we did for the generalized positions), the inner linear systems for Newton-type optimizers would possess a banded structure that enables more efficient direct solves without reordering of the system(s).
@@ 6.4 Toward More General Inverse Problems
In our framework, for simplicity and efficiency, we assumed a strand with constant radius, density, and material stiffness. It would be worthwhile removing this assumption to support more general anisotropic and inhomogeneous strands [HB21, HB23]. Moreover, our framework should be able to extend to support two-dimensional shells and three-dimensional volumetric structures, along with more rigorous analysis for the box constraints imposed on the rest shape parameters. In addition, although we also assumed no contacts among strands or other objects, this approach can lead to larger rest shape changes than necessary in cases where the elastic strands could be supported by each other or other objects when the simulation starts [DJBDDT13, LCBD18, HTYW22, HWP23]. Thus, a promising extension would be to take frictional contacts into account during the rest shape optimization.
Acknowledgements
We thank the anonymous reviewers for their valuable suggestions and comments, which significantly improved this work. We are also grateful to Bo Yang, Qingyuan Zheng, Rundong Wu, and Yu Ju Chen for their discussions during the early stages of this research. Additionally, we acknowledge the authors of [HMLL15] for publicly releasing the hair dataset. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2021-02524).